# Learning Unknown Intervention Targets in Structural Causal Models from Heterogeneous Data

### 11 Dec 2023 v1

@@ 4 Experiments
We evaluated the performance of LIT algorithm on randomly generated models. We considered the following three settings, with different numbers of environments $D=\{8,16,32\}$ for data generation: (1) Linear Gaussian model under causal sufficiency assumption; (2) Nonlinear model under causal sufficiency assumption; (3) Linear Gaussian model in the presence of latent confounders. We considered the following methods in our empirical studies: 1- LIT (our proposed method); 2- PreDITEr algorithm (Varici et al.,, 2022) which allows for latent confounders (that are not in ${\bf T}$) but assumes the model to be linear; 3- UT-IGSP algorithm (Squires et al.,, 2020) which works for both linear and nonlinear SCMs under merely causal sufficiency; 4- FCI-JCI123 algorithm in (Mooij et al.,, 2020) which allows for both latent confounders and nonlinearity in the model.
We repeated each setting for 40 times, and reported the average F1-score in recovering ${\bf T}_{{\bf O}}$ for each setting. The results are shown in Figure 2. (Full experimental results can be found in Appendix E.) Note that FCI-JCI is executable only under the first two settings with $D=8$ due to huge run times. The results of other algorithms for $D=8$ are provided in Appendix E. In the first setting, as we expected, PreDITEr has the best performance as it is designed specifically for linear Gaussian SCMs. UT-IGSP and LIT algorithm have decent performances, while FCI-JCI123 does not perform well. In the second setting, LIT and UT-IGSP can both recover the intervention targets with high accuracy. On the contrary, the performance of PreDITEr becomes worse as it is not designed for nonlinear models. Finally, in the third setting, in the presence of latent confounders, the performance of UT-IGSP becomes much worse because it cannot handle any latent confounders. Meanwhile, LIT outperforms PreDITEr and UT-IGSP for various numbers of variables in the system. Note that the best F1-score that can be achieved by any algorithm is strictly less than one as there are intervention targets that can never be recovered. Lastly, we note that LIT algorithm significantly reduces the number of CI tests performed: LIT algorithm takes at most 80 CI tests while PreDITER requires up to 30000 PDE estimates. See Appendix E for more details.
@@ 5 Comparison with previous work
Learning causal structures in multi-environment setting with unknown target interventions has been the focus of several recent work. Under causal sufficiency assumption, Ghassami et al., (2017) considered linear SCMs in multi-environment setting and showed that the intervention targets can be recovered by checking whether the variances of residuals of some linear regressions are changing across the environments. Squires et al., (2020) proposed an algorithm to recover an interventional Markov equivalence class (I-MEC) from a collection of interventional distributions with unknown intervention targets. The proposed algorithm greedily searches over the space of permutations to minimize a score function. In computing the score function, it is required to perform invariance tests to check whether pairs of distributions are equal which is in general sample inefficient in practice. Brouillard et al., (2020) proposed a causal discovery algorithm that recovers the intervention targets by finding the optimal solution of an appropriately defined score function. In both (Squires et al.,, 2020; Brouillard et al.,, 2020), no interventions are allowed in the observational environment and therefore the algorithm must know which environment pertains to the observation-only setting. More recently, Perry et al., (2022) proposed a score-based approach based on sparse mechanism shift hypothesis and showed that it can recover the true causal graph up to an I-MEC. However, it has the same drawback of performing invariance tests as in (Squires et al.,, 2020).
@@
All the aforementioned work learns the causal structure up to an equivalence class and identifies a candidate intervention set as a byproduct of the output of the causal discovery task. Very recently, Varici et al., (2022) proposed a method to learn merely intervention targets (rather than recovering the causal structure) in linear SCMs and it is more scalable than previous work as it does not perform extensive CI tests. Moreover, the identifiability result in recovering intervention targets is exactly the same as (Jaber et al.,, 2020). Note that all the previous work (Jaber et al.,, 2020; Mooij et al.,, 2020; Varici et al.,, 2022) assume that if latent confounders exist, none of them is in the intervention target set. We relax this assumption and allow latent confounders to be among the intervention targets.
@@ 6 Conclusions
We addressed the problem of identifying unknown intervention targets in a multi-environment setting. Our two-phase algorithm recovers the exogenous noises and matches them with corresponding endogenous variables. Under the causal sufficiency assumption, our algorithm uniquely identifies the intervention targets. In the presence of latent confounders, we provided a candidate intervention target set which is more informative than previous work. Experiment results support the advantages of the proposed algorithm in identifying intervention targets. As a future work, in the recovery phase, it is an interesting direction to strengthen the identifiability result of non-linear SCM with latent variables which would broaden the applicability of our method to more complex systems.
@@ A.1 Detailed description about contrastive learning approach and nonlinear ICA
How constrastive learning approach above is applied in our work
As we have access to a collection of distributions in $\mathcal{P}$, we will exploit the heterogeneity in data to recover the exogenous noises in ${\bf N}_{{\bf T}}$. Let auxiliary variable $U$ denote the index of the environment and assume that the exogenous noises belong to the exponential family in Definition 2. Moreover, assume that $\lambda_{i}$s corresponding to any $N_{i}\in{\bf N}_{{\bf T}}$ are randomly generated across the environments and $\tilde{q}_{i}(N_{i})$s are strictly monotonic functions of $|N_{i}|$. We utilize a contrastive learning approach similar to what we discussed above. We train the nonlinear regression model $r({\bf O},U)={\bf h}({\bf O})^{T}{\bf v}(U)+a({\bf O})+b(U)$ with universal approximation capability for the supervised learning task of discriminating $({\bf O},U)$ from $({\bf O},U^{\prime})$ where ${\bf h}({\bf O}):\mathbb{R}^{|{\bf O}|}\rightarrow\mathbb{R}^{|{\bf O}|}$, ${\bf v}(U):\mathbb{R}\rightarrow\mathbb{R}^{|{\bf O}|}$, and $a({\bf O})$, $b(U)$ are some scalar-valued functions.
@@ A.2 Additional discussion on Assumption 1
Second, for nonlinear SCMs, invertibility of the mixing function is satisfied when the model is an additive noise model, which can be written as $X_{i}=f_{i}(Pa_{i})+N_{i}$. In this case, the inverse function that maps ${\bf N}$ to ${\bf O}$ can be constructed according to the following equation: $N_{i}=X_{i}-f_{i}(Pa_{i})$. Note that the invertibility of the mixing function does not depend on $f_{i}$. This result can be generalized into the following remark:
@@
We note that the condition in Remark 3 can be satisfied when $f_{i}$ in the data generating model is designed as a multi-layer perceptron with ReLU activation function, where all model weights are positive.
@@ Appendix B Further discussion on the matching phase
B.1 Markov Property in the Augmented Graph
@@
B.2 Example Explaining the Conditions under Causal Sufficiency Assumption
@@
Figure 3 depicts the augmented graph of an SCM in which ${\bf T}=\{X_{1},X_{2},X_{5}\}$ (indicated by red circles). In the recovery phase, we recover three noises $\tilde{N}_{1},\tilde{N}_{2},\tilde{N}_{5}$, which are invertible transformations of $N_{1}$, $N_{2}$, $N_{5}$, respectively. Note that we do not know the correspondence of the noises to the variables as there are permutation indeterminacy. The indicator sets for all the variables are: ${\bf I}_{1}=\{\tilde{N}_{1}\}$, ${\bf I}_{2}=\{\tilde{N}_{2}\}$, ${\bf I}_{3}=\{\tilde{N}_{1},\tilde{N}_{2}\}$, ${\bf I}_{4}=\emptyset$, ${\bf I}_{5}={\bf I}_{6}={\bf I}_{7}=\{\tilde{N}_{1},\tilde{N}_{2},\tilde{N}_{5}\}$. For $X_{1}$ and $X_{2}$, the condition in II is satisfied. Thus, they are in ${\bf T}$. As for $X_{3}$ and $X_{4}$, the condition in I holds and therefore they are not in ${\bf T}$. For the variables in $\{X_{5},X_{6},X_{7}\}$, the condition in 5 holds and the only variable satisfying the condition in (C1) is $X_{5}$ as $\tilde{N}_{5}$ is independent of $X_{6}$ and $X_{7}$ given ${\bf S}_{5}\cup\{X_{5}\}$ where ${\bf S}_{5}=\{X_{1},X_{2},X_{3},X_{4}\}$.
B.3 Definition of Auxiliary Graph
In this section we present the formal definition of auxiliary graph.
Definition 1 (Auxiliary graph).
For each variable $X_{i}$, denote ${\bf I}_{0}(X_{i})$ as $An_{G_{\bf T}}(X_{i})\cap{\bf N}_{\bf T}$. Given a SCM and its corresponding augmented graph $G_{\bf T}$, the auxiliary graph $Aux(G_{\bf T})$ is constructed from $G_{\bf T}$ as follows:
  1. (a)
  2. (b)
    (i) For each $N_{l}\in{\bf N}_{\bf T}$ (noise that corresponds to a variable in ${\bf T_{L}}$) and each of its child $X_{i}$, keep the edge $N_{l}\rightarrow X_{i}$ if for any other child $X_{j}$ of $N_{l}$ in $G_{\bf T}$, ${\bf I}_{0}(X_{i})\subseteq{\bf I}_{0}(X_{j})$. Otherwise remove the edge $N_{l}\rightarrow X_{i}$. (ii) For each remaining edge $N_{l}\rightarrow X_{i}$, add (remove) the edge $N_{l}\rightarrow X_{k}$ if there is an (no) inducing path between $X_{k}$ and all (some) parents of $X_{i}$ in ${\bf N}_{\bf T}$ relative to ${\bf L}\setminus{\bf T_{L}}$ in $G_{\bf T}$, and ${\bf I}_{0}(X_{k})={\bf I}_{0}(X_{i})$.
Recall from Section 3.3.1 that under Assumption 2, $An_{G_{\bf T}}(X_{i})\cap{\bf N}_{\bf T}$ can be recovered by ${\bf I}_{i}$. Therefore we use ${\bf I}_{0}(X_{i})$ to represent the true value of ${\bf I}_{i}$ which does not depend on the recovery output. Further, it is noteworthy that while part (b)(ii) in Definition 1 includes both adding and removing of the edges, the addition or removal of edges in the step is independent of the order of the edge selection $N_{l}\rightarrow X_{i}$.
@@ D.1 Proof of Proposition 1
D.2 Regarding Assumption 1(b) on Linear non-Gaussian Models
@@ D.3 Proof of Proposition 2
Based on what we proved above, we can imply that ${\bf I}_{i}$ contains only the noises in $\tilde{{\bf N}}_{{\bf T}}$ whose corresponding variables are ancestors of $X_{i}$.
@@
II As the condition in I is not satisfied, there exists $k$ such that $\tilde{N}_{k}\in\mathcal{N}_{i}={\bf I}_{i}\backslash\cup_{j:X_{j}\in{\bf S}_{i}}{\bf I}_{j}$. Suppose that $\tilde{N}_{k}$ corresponds to $X_{k}$. Based on Remark 5, $X_{k}$ should be an ancestor of $X_{i}$. Moreover, we have ${\bf I}_{k}\subseteq{\bf I}_{i}$. As ${\bf I}_{i}$ is unique, then ${\bf I}_{k}\subsetneq{\bf I}_{i}$ and $X_{k}\in{\bf S}_{i}$. This contradictsthe fact that $\tilde{N}_{k}\in{\bf I}_{i}\backslash\cup_{j:X_{j}\in{\bf S}_{i}}{\bf I}_{j}$ and the proof is complete.
5 As the indicator sets do not satisfy the condition in I, $\mathcal{N}_{i}$ is not empty. Suppose that $\tilde{N}_{k}\in\mathcal{N}_{i}$ and $\tilde{N}_{k}$ corresponds to $X_{k}$. We know that $X_{k}$ should be in the set $\{X_{i_{1}},\cdots,X_{i_{p}}\}$. Otherwise, based on Remark 5, ${\bf I}_{k}\subseteq{\bf I}_{i_{1}}$. Now, if ${\bf I}_{k}\subsetneq{\bf I}_{i_{1}}$, then $X_{k}\in{\bf S}_{i_{1}}$ which is a contradiction. Thus, ${\bf I}_{k}$ should be equal to ${\bf I}_{i}$ but in that case $X_{k}$ is in the set $\{X_{i_{1}},\cdots,X_{i_{p}}\}$. Thus, we can conclude that at least one of the variables in this set is in ${\bf T}$. Please note that at most one of the variable in $\{X_{i_{1}},\cdots,X_{i_{p}}\}$ can be in ${\bf T}$. Otherwise, the variables in this set cannot have the same indicator set. Therefore, exactly one of the variables in $\{X_{i_{1}},\cdots,X_{i_{p}}\}$ is in ${\bf T}$ and we can obtain the corresponding recovered noise by ${\bf I}_{i_{1}}\backslash\cup_{j:X_{j}\in{\bf S}_{i_{1}}}{\bf I}_{j}$. Without loss of generality, suppose that $X_{i_{1}}\in{\bf T}$. Then, based on Remark 5, other variables in $\{X_{i_{1}},\cdots,X_{i_{p}}\}$ should be descendant of $X_{i_{1}}$. The set ${\bf S}_{i_{1}}$ contains the ancestors of $X_{i_{1}}$. Thus, it also includes parents of $X_{i_{1}}$. Now, for any path between $N_{i_{1}}$ and $X_{i_{j}}$, $j\geq 2$, if it is outgoing from node $X_{i_{1}}$, then it is blocked by $X_{i_{1}}$. If it is in-going toward $X_{i_{1}}$, then it is blocked by a parent of $X_{i_{1}}$ which is inside the set ${\bf S}_{i_{1}}$. Thus, under T-faithfulness assumption, it implies that $\tilde{N}_{i_{1}}\perp\!\!\!\perp X_{i_{j}}|\{X_{i_{1}}\}\cup{\bf S}_{i_{1}},\mbox{ for all }2\leq j\leq p$. Please note that the variables in $\{X_{i_{2}},\cdots,X_{i_{p}}\}$ cannot satisfy the condition in 5 as they cannot block the path $N_{i_{1}}\rightarrow X_{i_{1}}$ by any set ${\bf S}_{i_{j}}$ for $j\geq 2$ and the proof is complete.

### 09 Mar 2024 v2

@@ 4 Experiments
We evaluated the performance of LIT algorithm on randomly generated models. [^6] Our code is available at: https://github.com/Yuqin-Yang/LIT. Specifically, we considered the following three settings for data generation, and each with number of environments $D=\{8,16,32\}$: (1) Linear Gaussian model under causal sufficiency assumption; (2) Nonlinear model under causal sufficiency assumption; (3) Linear Gaussian model in the presence of latent confounders. We considered the following approaches in our empirical studies: 1- LIT (our proposed method); 2- PreDITEr algorithm (Varici et al.,, 2022), which allows for latent confounders (that are not in ${\bf T}$) but assumes the model is linear; 3- UT-IGSP algorithm (Squires et al.,, 2020), which works for both linear and nonlinear SCMs but under causal sufficiency; 4- CITE algorithm in (Varici et al.,, 2021), which only works for linear SCMs under causal sufficiency; 5- FCI-JCI123 algorithm in (Mooij et al.,, 2020), which allows for both latent confounders as well as nonlinearity in the model.
We repeated each setting for 40 times, and reported the average F1-score in recovering ${\bf T}_{{\bf O}}$ for each setting. The results are shown in Figure 3. (Full experimental results can be found in Appendix E.) Note that FCI-JCI is executable only under the first two settings with $D=8$ due to huge run times. The results of other algorithms for $D=8$ and $D=16$ are provided in Appendix E. In the first setting, as we expected, PreDITEr and CITE have the best performances as they are designed specifically for linear Gaussian SCMs. UT-IGSP and LIT algorithm have decent performances, while FCI-JCI123 does not perform well. In the second setting, LIT and CITE can both recover the intervention targets with high accuracy. On the contrary, the performance of PreDITEr becomes worse as it is not designed for nonlinear models. Finally, in the third setting, in the presence of latent confounders, the performance of UT-IGSP and CITE becomes much worse because they cannot handle any latent confounders. Meanwhile, LIT outperforms baseline algorithms for various numbers of variables in the system. Note that the best F1-score that can be achieved by any algorithm is strictly less than one as there are observed variables that are theoretically indistinguishable from the true intervention targets. Lastly, we note that LIT algorithm significantly reduces the number of CI tests performed: LIT algorithm takes at most 80 CI tests while PreDITER requires up to 30000 PDE estimates. See Appendix E for more details.
@@ 5 Comparison with previous work
Learning causal structures in multi-environment setting with unknown target interventions has been the focus of several recent work. Under the causal sufficiency assumption, Ghassami et al., (2017) considered linear SCMs in multi-environment setting and showed that the intervention targets can be recovered by checking whether the variances of residuals pertaining to certain linear regressions change across the environments. Squires et al., (2020) proposed an algorithm to recover an interventional Markov equivalence class (I-MEC) from a collection of interventional distributions with unknown intervention targets. The proposed algorithm greedily searches over the space of permutations to minimize a score function. In computing the score function, it is required to perform invariance tests to check whether pairs of distributions are equal which is in general sample inefficient in practice. Brouillard et al., (2020) proposed a causal discovery algorithm that recovers the intervention targets by finding the optimal solution of an appropriately defined score function. In both (Squires et al.,, 2020; Brouillard et al.,, 2020), no interventions are allowed in the observational environment and therefore the algorithm must know which environment pertains to the observation-only setting. More recently, Perry et al., (2022) proposed a score-based approach based on sparse mechanism shift hypothesis and showed that it can recover the true causal graph up to an I-MEC. However, it has the same drawback of performing invariance tests as in (Squires et al.,, 2020).
@@
All the aforementioned work learns the causal structure up to an equivalence class and identifies a candidate intervention set as a byproduct of the output of the causal discovery task. Very recently, Varici et al., (2021, 2022) proposed two methods to learn merely intervention targets (rather than recovering the causal structure) in linear Gaussian SCMs with or without latent confounders. Their methods are more scalable than previous work as they do not perform extensive CI tests. Moreover, the identifiability result in recovering intervention targets is exactly the same as (Jaber et al.,, 2020). Note that all the previous work (Jaber et al.,, 2020; Mooij et al.,, 2020; Varici et al.,, 2022) assume that if latent confounders exist, none of them is in the intervention target set. We relax this assumption and allow latent confounders to be among the intervention targets.
@@ 6 Conclusions
We addressed the problem of identifying unknown intervention targets in a multi-environment setting. In particular, we considered the case when latent confounders are allowed to be in the intervention target set. This may happen when we can only partially observe the system variables. For instance, this is the case in genomics data, where it might be the case that not all affected genes or proteins are measured (Verny et al.,, 2017). Existing approaches such as $\psi$-FCI and PreDITEr do not allow for changing latent confounders and lack theoretical guarantees in such scenarios. Our two-phase algorithm recovers the exogenous noises and matches them with corresponding endogenous variables. Under the causal sufficiency assumption, our algorithm uniquely identifies the intervention targets. In the presence of latent confounders, we provided a candidate intervention target set which is more informative than previous work. Experiment results support the advantages of the proposed algorithm in identifying intervention targets. Our simulations showed existing methods might misinterpret changes in observed variables influenced by these latent confounders, leading to lower precision. In contrast, our method effectively handles latent confounders in intervention targets. As a future work, in the recovery phase, it is an interesting direction to strengthen the identifiability result of non-linear SCM with latent variables which would broaden the applicability of our method to more complex systems.
Acknowledgements
Negar Kiyavash’s research was in part supported by the Swiss National Science Foundation under NCCR Automation, grant agreement 51NF40_180545 and Swiss SNF project 200021_204355 /1.
@@ A.1 Detailed Description about Contrastive Learning Approach and Nonlinear ICA
How constrastive learning approach above is applied in our work.
As we have access to a collection of distributions in $\mathcal{P}$, we will exploit the heterogeneity in data to recover the exogenous noises in ${\bf N}_{{\bf T}}$. Let auxiliary variable $U$ denote the index of the environment and assume that the exogenous noises belong to the exponential family in Definition 2. Moreover, assume that $\lambda_{i}$’s corresponding to any $N_{i}\in{\bf N}_{{\bf T}}$ are randomly generated across the environments and $\tilde{q}_{i}(N_{i})$s are strictly monotonic functions of $|N_{i}|$. We utilize a contrastive learning approach similar to what we discussed above. We train the nonlinear regression model $r({\bf O},U)={\bf h}({\bf O})^{T}{\bf v}(U)+a({\bf O})+b(U)$ with universal approximation capability for the supervised learning task of discriminating $({\bf O},U)$ from $({\bf O},U^{\prime})$ where ${\bf h}({\bf O}):\mathbb{R}^{|{\bf O}|}\rightarrow\mathbb{R}^{|{\bf O}|}$, ${\bf v}(U):\mathbb{R}\rightarrow\mathbb{R}^{|{\bf O}|}$, and $a({\bf O})$, $b(U)$ are some scalar-valued functions.
@@ A.2 Additional Discussion on Assumption 1
Second, for nonlinear SCMs, the invertibility of the mixing function is satisfied when the model is an additive noise model, which can be written as $X_{i}=f_{i}(Pa_{i})+N_{i}$. In this case, the inverse function that maps ${\bf N}$ to ${\bf O}$ can be constructed according to the following equation: $N_{i}=X_{i}-f_{i}(Pa_{i})$. Note that the invertibility of the mixing function does not depend on $f_{i}$. This result can be generalized into the following remark:
@@
We note that the condition in Remark 3 can be satisfied when $f_{i}$ in the data generating model is designed as a multi-layer perceptron with ReLU activation function, where all model weights are strictly positive.
@@ Appendix B Further discussion on the matching phase
B.1 Markov Property in the Augmented Graph $G_{{\bf T}}$
@@
[…]
B.2 Example Explaining the Conditions under the Causal Sufficiency Assumption
@@
Figure 4 depicts the augmented graph of an SCM in which ${\bf T}=\{X_{1},X_{2},X_{5}\}$ (indicated by red circles). In the recovery phase, we recover three noises $\tilde{N}_{1},\tilde{N}_{2},\tilde{N}_{5}$, which are invertible translations of $N_{1}$, $N_{2}$, $N_{5}$, respectively. Note that we do not know the correspondence of the noises to the variables as there are permutation indeterminacy. The indicator sets for all the variables are: ${\bf I}_{1}=\{\tilde{N}_{1}\}$, ${\bf I}_{2}=\{\tilde{N}_{2}\}$, ${\bf I}_{3}=\{\tilde{N}_{1},\tilde{N}_{2}\}$, ${\bf I}_{4}=\emptyset$, ${\bf I}_{5}={\bf I}_{6}={\bf I}_{7}=\{\tilde{N}_{1},\tilde{N}_{2},\tilde{N}_{5}\}$. For $X_{1}$ and $X_{2}$, the condition in II is satisfied. Thus, they are in ${\bf T}$. As for $X_{3}$ and $X_{4}$, the condition in I holds and therefore they are not in ${\bf T}$. For the variables in $\{X_{5},X_{6},X_{7}\}$, the condition in 5 holds and the only variable satisfying the condition in (C1) is $X_{5}$ as $\tilde{N}_{5}$ is independent of $X_{6}$ and $X_{7}$ given ${\bf S}_{5}\cup\{X_{5}\}$ where ${\bf S}_{5}=\{X_{1},X_{2},X_{3},X_{4}\}$.
@@ D.1 Proof of Proposition 1
D.2 Regarding Assumption 1(b) in Linear non-Gaussian Models
@@ D.3 Proof of Proposition 2
Based on what we proved above, we can imply that ${\bf I}_{i}$ contains only the noises in $\tilde{{\bf N}}_{{\bf T}}$ whose corresponding variables are ancestors of $X_{i}$ in $G_{\bf T}$.
@@
II As the condition in I is not satisfied, there exists $\tilde{N}_{k^{\prime}}\in\mathcal{N}_{i}={\bf I}_{i}\backslash\cup_{j:X_{j}\in{\bf S}_{i}}{\bf I}_{j}$. Suppose that $\tilde{N}_{k^{\prime}}$ corresponds to $X_{k}$. Based on Remark 7, $X_{k}$ should be an ancestor of $X_{i}$. Moreover, we have ${\bf I}_{k}\subseteq{\bf I}_{i}$. As ${\bf I}_{i}$ is unique, then ${\bf I}_{k}\subsetneq{\bf I}_{i}$ and $X_{k}\in{\bf S}_{i}$. But this is in contradiction with the assumption $k^{\prime}\in{\bf I}_{i}\backslash\cup_{j:X_{j}\in{\bf S}_{i}}{\bf I}_{j}$ and the proof is complete.
5 As the indicator sets do not satisfy the condition in I, $\mathcal{N}_{i}={\bf I}_{i_{1}}\backslash\cup_{j\in{\bf S}_{i_{1}}}{\bf I}_{j}$ is not empty. Suppose that $\tilde{N}_{k^{\prime}}\in\mathcal{N}_{i}$ and $\tilde{N}_{k^{\prime}}$ corresponds to $X_{k}$. We know that $X_{k}$ should be in the set $\mathcal{X}_{i}=\{X_{i_{1}},\cdots,X_{i_{p}}\}$. Otherwise, based on Remark 7, ${\bf I}_{k}\subseteq{\bf I}_{i_{1}}$. Now, if ${\bf I}_{k}\subsetneq{\bf I}_{i_{1}}$, then $X_{k}\in{\bf S}_{i_{1}}$ which is a contradiction. Thus, ${\bf I}_{k}$ should be equal to ${\bf I}_{i}$ but in that case $X_{k}$ is in the set $\mathcal{X}_{i}$. Thus, we can conclude that at least one of the variables in this set is in ${\bf T}$. Please note that at most one of the variable in $\mathcal{X}_{i}$ can be in ${\bf T}$. Otherwise, the variables in this set cannot have the same indicator set. Therefore, exactly one of the variables in $\mathcal{X}_{i}$ is in ${\bf T}$ and we can obtain the corresponding recovered noise, which is the only recovered noise in $\mathcal{N}_{i}$. Without loss of generality, suppose that $X_{i_{1}}\in{\bf T}$. Then, based on Remark 7, other variables in $\{X_{i_{2}},\cdots,X_{i_{p}}\}$ should be descendant of $X_{i_{1}}$. The set ${\bf S}_{i_{1}}$ contains the ancestors of $X_{i_{1}}$. Thus, it also includes parents of $X_{i_{1}}$. Now, for any path between $N_{i_{1}}$ and $X_{i_{j}}$ in the augmented graph $G_{\bf T}$, $j\geq 2$, if it is outgoing from node $X_{i_{1}}$, then it is blocked by $X_{i_{1}}$. If it is in-going toward $X_{i_{1}}$, then it is blocked by a parent of $X_{i_{1}}$ which is inside the set ${\bf S}_{i_{1}}$. Thus, we can imply that $\tilde{N}_{i_{1}}\perp\!\!\!\perp X_{i_{j}}|\{X_{i_{1}}\}\cup{\bf S}_{i_{1}},\mbox{ for all }2\leq j\leq p$ under Assumption 2. Please note that the variables in $\{X_{i_{2}},\cdots,X_{i_{p}}\}$ cannot satisfy the condition in 5 as they cannot block the path $N_{i_{1}}\rightarrow X_{i_{1}}$ by any set ${\bf S}_{i_{j}}$ for $j\geq 2$ and the proof is complete.